OpenStudy (anonymous):

i am having trouble with this problem -2 x (-2)-(2-(-3) exponent 2) exponent 2

OpenStudy (anonymous):

-2*(-2)-(2-((-3)^2)))^2 = -2*(-2)-(2-9)^2 = -2*(-2)-49 = -45 (-3)^2 = 9 (2-9)^2 = (-7)^2 = 49 -2*(-2) = 4 4-49 = 45